Question 2 A sample of 6.2 g of copper (II) carbonate is heated until a constant mass is obtained. Calculate the mass of copper(II) oxide that will be formed from 6.2 g of copper(II) carbonate. (Relative atomic mass of C:12, 0:16, Cu: 64) (10 marks]


1
Expert's answer
2021-07-19T04:05:15-0400

Molar Mass of CuCO3 = 124

= 6.2/124= 0.05

Molar Mass of copper (II) Oxide = 80

= 80×0.05

= 4g